# Training on the Test Task

### v2

@@ Abstract
We study a fundamental problem in the evaluation of large language models that we call training on the test task. Unlike wrongful practices like training on the test data, leakage, or data contamination, training on the test task is not a malpractice. Rather, the term describes a growing set of practices that utilize knowledge about evaluation tasks at training time. We demonstrate that training on the test task confounds both relative model evaluations and claims about emergent capabilities. We argue that the seeming superiority of one model family over another may be explained by a different degree of training on the test task. To this end, we propose an effective method to adjust for the effect of training on the test task on benchmark evaluations. Put simply, to fine-tune each model under comparison on the same task-relevant data before evaluation. We then show that instances of emergent behavior disappear gradually as models train on the test task. Our work promotes a new perspective on the evaluation of large language models with broad implications for benchmarking and the study of emergent capabilities. [^2] Code and fine-tuned models are available at https://github.com/socialfoundations/training-on-the-test-task
@@ 1 Introduction
But it’s been much less clear what to do about legitimate attempts to bring training closer to evaluation. There is an obvious a gap between next token prediction at training time and tasks, such as reasoning and question answering, at test time. Ongoing research and engineering efforts, in fact, aim to narrow precisely this gap (MetaAI, 2024; Lewis, 2024). Why shouldn’t training be informed by knowledge about the downstream test tasks? What’s an unfair advantage of some may be the feature of others.
In this work, we group strategies to utilize task knowledge at training time under the umbrella term of training on the test task. Examples of training on the test task include the use of instruction-tuning data or question answering templates during pre-training (Bai et al., 2023; StabilityAI, 2023; Groeneveld et al., 2024; Zhang et al., 2024). Models may also implicitly train on the test task when their pretraining data is selected through ablations on downstream benchmark evaluations (Gemma et al., 2024; MetaAI, 2024). We work from the premise that training on the test task is acceptable–or at least, unavoidable.
@@ 1.1 Our contributions
We propose a simple and effective method to adjust for the effect of training on the test task on benchmark evaluations. Put simply, to fine-tune each model on the same, sufficient amount of task-specific data before evaluation. To validate our method, we demonstrate its effectiveness in a controlled setting: we take the older models and fine-tune them on the test task. Remarkably this recreates the kind of performance differences observed between newer and older models, further suggesting that training on the test task explains the improvements of newer models. We then show that we can undo the advantage of the fine-tuned models over the other models by further fine-tuning all models on the test task (Section 3.1, Figure 3).
Next, we give evidence that training on the test task may be a more dominant factor in benchmark performance than data contamination. To argue this point, we consider ARC and HellaSwag. Here, at first there appears to be no sign of newer models have an unfair advantage over older models. But after reformulating these benchmarks as MMLU-style multiple choice question answering tasks (MCQA), we see the same confounded results as for MMLU (Section 3.2, Figure 4). This suggest that the improvements of newer models on MMLU are likely not because of memorization of specific testing data, but rather due to an improved ability for MCQA tasks. Either way, our proposed adjustment recovers fair model comparisons.
Then, we show how training on the test task distorts model family comparisons. Certain model families appear markedly superior to others before adjusting for test task training but not after adjustment (Section 4.1, Figure 6). We then demonstrate how training on the test task has inflated the perceived progress made by recent model families. After adjusting for its effect, newer models offer only modest improvements to the Pareto frontier of model performance relative to pre-training compute (Section 4.2, Figure 7).
@@
Our work calls for a major reorientation of large language model evaluation. Model comparisons and claims of emergence are strongly confounded by the choice of training data relative to the test tasks. Rather than scrambling to detect and disallow various forms of training on the test task, we propose to “fight fire with fire”. When comparing models with different pre-training data, our recommendation is to give each model the same sufficient amount of fine-tuning on task-relevant data prior to evaluation.
@@ 2 Adjusting for training on the test task
We choose MMLU (Hendrycks et al., 2020) and GSM8K (Cobbe et al., 2021) as a case study for investigating training on the test task in active benchmarks. MMLU tests for world knowledge, whereas GSM8K tests multi-step mathematical reasoning. These two benchmarks are arguably the most influential of the 2022-2024 period under study [^1] For example, GPT 4 (Achiam et al., 2023), Claude 3 (Anthropic, 2024), Gemini (Gemini et al., 2023) and Llama 3 (MetaAI, 2024) all report and prominiently highlight MMLU and GSM8K performance. They are also included in the HuggingFace (HF) Open LLM Leaderboard v1 (Beeching et al., 2023), a popular leaderboard that evaluates and ranks models with publicly available weights. We evaluate models using LM Evaluation Harness (EleutherAI, 2024), in identical fashion to the HF leaderboard [^2] Similarly to the HF leaderboard, we evaluate MMLU and GSM8K 5-shot, ARC 25-shot, and HellaSwag 10-shot..
@@
We evaluate models on MMLU and GSM8K, and plot benchmark accuracy against pretraining compute in Figure 1 top. We observe that performance correlates with pretraining compute for both benchmarks. However, on the surface it appears that models released after November 2023 better leverage pretraining compute. In other words, for a given compute budget newer models are able to attain better benchmark performance.
We selected November 2023 as the temporal cutoff for our analysis because the technical reports of models released from late 2023 onward start referencing certain pre-training practices that may amount to training on test task. For example, Qwen (Bai et al., 2023), Olmo 1.7 (Groeneveld et al., 2024) and MAP Neo (Zhang et al., 2024) explicitly include instruction data during pretraining. StableLM 2 (StabilityAI, 2023) reformulates some of its pretraining datasets to better resemble downstream tasks such as question-answering. More subtly, the pretraining data mixtures of Gemma (Gemma et al., 2024) and Llama 3 (MetaAI, 2024) were determined through extensive ablations on downstream benchmark evaluations. We validate that our findings are robust to adjusting the temporal cutoff by a few months; see Appendix C.1 for details. Choosing specifically the month of November as the cutoff is therefore not critical for our analysis.
@@ 2.1 Adjusting for training on the test task by training on the test task
Remarkably, we observe that older models tend to benefit much more from fine-tuning on task-relevant data compared to newer models, see Figure 2. The improvements in older models are striking, often leaping from random chance accuracy to double-digit gains in accuracy. In contrast, fine-tuning provides comparatively little benefit to newer models. This observation suggests that newer models have already been exposed to a substantial amount of task-relevant data, making additional fine-tuning less impactful.
@@ 2.2 Quantifying performance differences between newer and older models
Before adjusting for test task training, the estimated difference in performance $\widehat{\theta}$ between newer and older models are statistically significant, positive, and large. Specifically, recent models outperform older ones on average by over 7 accuracy points in MMLU and 19 accuracy points in GSM8K. These are remarkably large differences in benchmark performance. However, after the adjustment, the estimated coefficient $\widehat{\theta}$ is both small and not statistically significant. See Figure 1 bottom. That is, conditioned on all models training on the same amount of task-specific data, we find no evidence for a significant difference in benchmark performance between newer and older models.
@@ 3.1 Fine-tuning on the test task
For this section, we only consider models trained before November 2023. We split models into two cohorts: a control group and a treatment group. We take these older models as a control group. We then create a treatment group by fine-tuning the control group on the datasets of task-relevant data introduced in Section 2. We only fine-tune models with at least $7\cdot 10^{21}$ FLOPs, the pre-training compute of the smallest newer model. We fine-tune for a single epoch. We plot in Figure 3 top the benchmark performance of the two cohorts.
Figure 3: Models trained before November 2023 () without fine-tuning and () after fine-tuning on the test task. Their difference in benchmark performance $\widehat{\theta}$ resemble that between newer and older models. After adjusting by training on the test task, their difference vanishes. Bold indicates significance with $p<0.05$.
@@ 3.2 Reformulating the test task
In this section we show that reformulating other benchmarks as multiple-choice question answering tasks leads to similar differences in performance between older and newer models. We consider two additional benchmarks from the HF leaderboard v1: ARC Challenge (Clark et al., 2018) and HellaSwag (Zellers et al., 2019). Similarly to MMLU, ARC is comprised of grade-school level questions. HellaSwag instead tests for commonsense reasoning. Like MMLU, the questions in ARC and HellaSwag are accompanied by four possible answers, among which the model must differentiate the correct one. ARC and HellaSwag use “cloze” evaluations: a models’ answer is taken to be that with the largest completion likelihood given the input question. In contrast, MMLU formulates questions as multiple-choice: all four answer choices are listed, and the model is promoted to pick one of the answer choices.
@@ 4 Implications for model comparisons
So far, we have shown how training on the test task distorts benchmark evaluations. Next, we examine its impact on the relative comparison of model families (Section 4.1), as well as its implications for accurately measuring progress in model capabilities over time (Section 4.2).
@@ 4.1 Comparing model families
Training on the test task therefore profoundly confounds relative model comparisons. Base models are rarely used “as is” and are generally adapted before deployment. Because of the confounding of training on the test task, performance before adaptation may not reliably predict performance after adaptation.
Interestingly, recent work equates pretraining data quality with downstream benchmark performance (Penedo et al., 2024; Li et al., 2024). Our findings suggest that a major contributing factor to the superior performance of “higher quality” pretraining datasets may be that they simply contain more test task data.
@@ 4.2 Progress in model capabilities
On the other hand, recent models tend to be trained on more data than Chinchilla compute-optimal (Hoffmann et al., 2022). Given the Chinchilla scaling laws, it is noteworthy that newer, smaller “over-trained” models match the performance of older, larger ones for the same amount of pretraining compute.
@@ 5 Implications for emergence
Throughout our evaluations, we observe emergent behaviour for MMLU and GSM8K: models perform at near random chance up to a certain scale of pretraining compute, followed by relatively sharper improvements in performance at larger scales (Wei et al., 2022b). After training on the test task, however, emergence for MMLU and GSM8K appears to occur at substantially lower scales. We dedicate this section to more closely investigate the relationship between training on the test task and emergence.
@@
We consider only models trained before November 2023, as we have established that these models train on the test task less than newer models. We evaluate the models at intermediate checkpoints as we fine-tune them on the datasets of task relevant data introduced in Section 2.1. We fit $\alpha$ and $c_{e}$ in Equation 1 to the different intermediate checkpoints, and report in Figure 8 top the corresponding points of emergence $c_{e}$. We find that emergence arises at increasingly lower compute regimes as models train on the test task. For MMLU, models exhibit emergence at around $10^{22}$ FLOPs, the scale of Pythia 6.9B. After training on 64,000 examples, emergence arises around around $6\cdot 10^{20}$ FLOPs, the scale of Pythia 410M. We observe similar results for GSM8K, see Figure 18 in Appendix E.
Figure 8: Scaling on MMLU as models increasingly train on the test task. The point of emergence $c_{e}$ arises at lower scales (top). Training on the test task yields claner log-linear scaling fits (bottom).
@@
The log-linear relationship between pretraining loss and compute is well-established (Kaplan et al., 2020). We observe that, for the compute ranges that we consider, training on the test task increasingly recovers log-linear scaling between pretraining compute and benchmark accuracy. Similarly to the earlier section, we evaluate intermediate checkpoints but instead fit log-linear functions in Figure 8 bottom. We observe that the R2 of the fit improves substantially as the models train on more task-relevant data, jumping from $0.63$ to $0.95$ after training on 64,000 examples. Therefore, after training on the test task almost all of the variation in benchmark accuracy is explained by log-linear scaling of pre-training compute. We observe similar results for GSM8K, see Figure 18 in Appendix E.
@@
Schaeffer et al. (2024a) argue that emergence appears due to the choice of metric. To mitigate emergence, they suggest to consider Brier score instead of accuracy. We observe, however, that emergence for MMLU does not disappear when using the Brier score, see Figure 5 right, nor that of ARC and HellaSwag when framed as multiple-choice questions, see Figure 17 in Appendix E. We discuss two practical solutions to obtain predictive scaling while maintaining accuracy as the evaluation metric.
For MMLU and multiple-choice benchmarks more broadly, cloze evaluationsc consistently yield smoother and more predictable scaling even when using accuracy as the evaluation metric. Since the purpose of these benchmarks is knowledge-testing more so than testing multiple-choice answering ability, cloze evaluations are preferable insofar predictive scaling at lower compute scales is an important consideration. This recommendation aligns with the concurrent work by Gu et al. (2024).
More broadly, if sufficient task relevant data is available, then training on the test task can result in much more predictable scaling by shifting emergence to smaller compute scales. That is, scaling laws where models across scales are fine-tuned on the same, sufficient task-relevant data prior to evaluation. Such scaling laws correspond to those of specialist models, which for some tasks –e.g., legal annotation (Dominguez-Olmedo et al., 2024)– or purposes –e.g., safety– might be preferable to the scaling laws of generalist models.

### v3

@@ Abstract
We study a fundamental problem in the evaluation of large language models that we call training on the test task. Unlike wrongful practices like training on the test data, leakage, or data contamination, training on the test task is not a malpractice. Rather, the term describes a growing set of practices that utilize knowledge about evaluation tasks at training time. We demonstrate that training on the test task confounds both relative model evaluations and claims about emergent capabilities. We argue that the seeming superiority of one model family over another may be explained by a different degree of training on the test task. To this end, we propose an effective method to adjust for the effect of training on the test task on benchmark evaluations. Put simply, to fine-tune each model under comparison on the same task-relevant data prior to evaluation. We then show that instances of emergent behavior disappear gradually as models train on the test task. Our work promotes a new perspective on the evaluation of large language models, with broad implications for benchmarking and the study of emergent capabilities. [^2] Code and fine-tuned models are available at https://github.com/socialfoundations/training-on-the-test-task
@@ 1 Introduction
But it’s been much less clear what to do about legitimate attempts to bring training closer to evaluation. There is an obvious a gap between next token prediction at training time and tasks, such as reasoning and question answering, at test time. Ongoing research and engineering efforts, in fact, aim to narrow precisely this gap (MetaAI, 2024; Lewis, 2024). Why shouldn’t training be informed by knowledge about the downstream test tasks? What’s an unfair advantage for some may be the feature of others.
In this work, we group strategies to utilize task knowledge at training time under the umbrella term of training on the test task. Examples of training on the test task include the use of instruction-tuning data or question answering templates during pre-training (Bai et al., 2023; StabilityAI, 2023; Groeneveld et al., 2024; Zhang et al., 2024). Models may also implicitly train on the test task when their pretraining data is selected through ablations on downstream benchmark evaluations (Gemma et al., 2024; MetaAI, 2024). We work from the premise that training on the test task is acceptable–or, at least, unavoidable.
@@ 1.1 Our contributions
We propose a simple and effective method to adjust for the effect of training on the test task on benchmark evaluations. Put simply, to fine-tune each model on the same, sufficient amount of task-specific data before evaluation. To validate our method, we demonstrate its effectiveness in a controlled setting: we take the older models and fine-tune them on the test task. Remarkably, this recreates the kind of performance differences observed between newer and older models, further suggesting that training on the test task explains the improvements of newer models. We then show that we can undo the advantage of the fine-tuned models over the other models by further fine-tuning all models on the test task (Section 3.1, Figure 3).
Next, we provide evidence that training on the test task may be a more dominant factor in benchmark performance than data contamination. To argue this point, we consider ARC and HellaSwag. Here, at first, there appears to be no sign of newer models having an unfair advantage over older models. But after reformulating these benchmarks as MMLU-style multiple choice question answering tasks (MCQA), we see the same confounded results as for MMLU (Section 3.2, Figure 4). This suggests that the improvements of newer models on MMLU are likely not because of memorization of specific testing data, but rather due to an improved ability for MCQA tasks. Either way, our proposed adjustment recovers fair model comparisons.
Then, we show how training on the test task distorts model family comparisons. Certain model families appear markedly superior to others before adjusting for test task training, but not after adjustment (Section 4.1, Figure 6). Model rankings on MMLU and GSM8K significantly change after such ajustment (Section 4.2, Figure 7). We then demonstrate how training on the test task has inflated the perceived progress made by recent model families. After adjusting for its effect, newer models offer only modest improvements to the Pareto frontier of model performance relative to pre-training compute (Section 4.3, Figure 8).
@@
Our work calls for a major reorientation of large language model evaluation. Model comparisons and claims of emergence are strongly confounded by the choice of training data relative to the test tasks. Rather than scrambling to detect and disallow various forms of training on the test task, we propose to “fight fire with fire”. When comparing models with different pre-training data, our recommendation is to give each model the same sufficient amount of fine-tuning on task-relevant data before evaluation.
@@ 2 Adjusting for training on the test task
We choose MMLU (Hendrycks et al., 2020) and GSM8K (Cobbe et al., 2021) as a case study for investigating training on the test task in active benchmarks. MMLU tests for world knowledge, whereas GSM8K tests multistep mathematical reasoning. These two benchmarks are arguably the most influential of the 2022-2024 period under study [^1] For example, GPT 4 (Achiam et al., 2023), Claude 3 (Anthropic, 2024), Gemini (Gemini et al., 2023) and Llama 3 (MetaAI, 2024) all report and prominently highlight MMLU and GSM8K performance. They are also included in the HuggingFace (HF) Open LLM Leaderboard v1 (Beeching et al., 2023), a popular leaderboard that evaluates and ranks models with publicly available weights. We evaluate models using LM Evaluation Harness (EleutherAI, 2024), in identical fashion to the HF leaderboard [^2] Similarly to the HF leaderboard, we evaluate MMLU and GSM8K 5-shot, ARC 25-shot, and HellaSwag 10-shot..
@@
We evaluate models on MMLU and GSM8K, and plot benchmark accuracy against pretraining compute in Figure 1 top. We observe that performance correlates with pretraining compute for both benchmarks. However, on the surface, it appears that models released after November 2023 better leverage pretraining compute. In other words, for a given compute budget, newer models tend to attain better benchmark performance.
We selected November 2023 as the temporal cutoff for our analysis because the technical reports of models released from late 2023 onward start referencing certain pre-training practices that may amount to training on the test task. For example, Qwen (Bai et al., 2023), Olmo 1.7 (Groeneveld et al., 2024) and MAP Neo (Zhang et al., 2024) explicitly include instruction data during pretraining. StableLM 2 (StabilityAI, 2023) reformulates some of its pretraining datasets to better resemble downstream tasks such as question-answering. More subtly, the pretraining data mixtures of Gemma (Gemma et al., 2024) and Llama 3 (MetaAI, 2024) were determined through extensive ablations on downstream benchmark evaluations. We validate that our findings are robust to adjusting the temporal cutoff by a few months; see Appendix C.1 for details. Choosing specifically the month of November as the cutoff is therefore not critical for our analysis.
@@ 2.1 Adjusting for training on the test task by training on the test task
We observe that older models tend to benefit much more from fine-tuning on task-relevant data compared to newer models, see Figure 2. The improvements in older models are striking, often leaping from random chance accuracy to double-digit gains in accuracy. In contrast, fine-tuning provides comparatively little benefit to newer models. This observation suggests that newer models have already been exposed to a substantial amount of task-relevant data, making additional fine-tuning less impactful.
@@ 2.2 Quantifying performance differences between newer and older models
Before adjusting for test task training, the estimated difference in performance $\widehat{\theta}$ between newer and older models are statistically significant, positive, and large. Specifically, recent models outperform older ones on average by over 7 accuracy points in MMLU and 19 accuracy points in GSM8K. These are remarkably large differences in benchmark performance.
However, after the adjustment, the estimated coefficient $\widehat{\theta}$ is both small and not statistically significant. See Figure 1 bottom. That is, conditioned on all models training on the same amount of task-specific data, we find no evidence for a significant difference in benchmark performance between newer and older models.
@@ 3.1 Fine-tuning on the test task
For this section, we only consider models trained before November 2023. We split the models into two cohorts: a control group and a treatment group. We take these older models as a control group. We then create a treatment group by fine-tuning the control group on the datasets of task-relevant data introduced in Section 2. We only fine-tune models with at least $7\cdot 10^{21}$ FLOPs, the pre-training compute of the smallest newer model. We fine-tune for a single epoch. We plot in Figure 3 top the benchmark performance of the two cohorts, and repeat the statistical analysis of Section 2.2.
Figure 3: Models trained before November 2023 () without fine-tuning and () after fine-tuning on the test task. Their difference in benchmark performance $\widehat{\theta}$ resembles that between newer and older models. After adjusting by training on the test task, their difference vanishes. Bold indicates significance with $p<0.05$.
@@ 3.2 Reformulating the test task
In this section, we show that reformulating other benchmarks as multiple-choice question answering tasks leads to similar differences in performance between older and newer models. We consider two additional benchmarks from the HF leaderboard v1: ARC Challenge (Clark et al., 2018) and HellaSwag (Zellers et al., 2019). Similarly to MMLU, ARC comprises grade-school level questions. HellaSwag instead tests for commonsense reasoning. Like MMLU, the questions in ARC and HellaSwag are accompanied by four possible answers, among which the model must differentiate the correct one. ARC and HellaSwag use “cloze” evaluations: a models’ answer is taken to be that with the largest completion likelihood given the input question. In contrast, MMLU formulates questions as multiple-choice: all four answer choices are listed, and the model is promoted to pick one of the answer choices.
@@ 4 Implications for model comparisons
So far, we have shown how training on the test task distorts benchmark evaluations. Next, we examine its impact on the relative comparison of model families (Section 4.1) and model rankings (Section 4.2), as well as its implications for accurately measuring progress in model capabilities over time (Section 4.3).
@@ 4.1 Comparing model families
Training on the test task therefore profoundly confounds relative model comparisons. Base models are rarely used “as is” and are generally adapted before deployment. Because of the confounding of training on the test task, performance before post-training may not reliably predict performance after post-training. It therefore makes little sense to compare base models at face value.
Ruan et al. (2024) argue that model families only vary in their efficiency in converting training compute to capabilities. Our findings show that training on the test task is a highly effective–and possibly the primary–mechanism for improving such compute efficiency.
A different line of work equates pretraining data quality with downstream benchmark performance (Penedo et al., 2024; Li et al., 2024). Our findings suggest that a major contributing factor to the superior performance of “higher quality” pretraining datasets may be that they simply contain more test task data. For example, the top-performing dataset in the DataComp-LM competition (Li et al., 2024)DCLM-Baseline— uses a popular post-training dataset as a reference for filtering the pretraining data.
Figure 7: Training on the test task significantly alters model rankings on GSM8K.
4.2 Model rankings
Model rankings are a crucial scientific export of machine learning benchmarks (Hardt, 2024). Figure 7 shows the GSM8K rankings of the 56 models we analyzed, both before and after adjusting for training on the test task. After the adjustment, the average ranking shift in GSM8K is 7.7 ranks, with a maximum shift of 21 ranks. Similarly, MMLU rankings exhibit substantial changes, with an average shift of 4.8 ranks and a maximum of 16 (see Figure 18 in Appendix E). These findings indicate that beyond affecting absolute benchmark performance, training on the test task also significantly alters model rankings.
4.3 Progress in model capabilities
@@
On the other hand, recent models tend to be trained on more data than Chinchilla compute-optimal (Hoffmann et al., 2022). Given the Chinchilla scaling laws, it is noteworthy that newer, smaller “over-trained” models match the performance of older, larger ones for the same amount of pretraining compute. Since inference and fine-tuning of smaller models is substantially cheaper, recent models can be much more accessible to less well-resourced institutions, with little cost in performance. For example, we find that Llama 3 8B closely matches the performance of Llama 2 70B (both have similar pre-training compute).
@@ 5 Implications for emergence
Throughout our evaluations, we observe emergent behavior for MMLU and GSM8K: models perform at near random chance up to a certain scale of pretraining compute, followed by relatively sharper improvements in performance at larger scales (Wei et al., 2022b). After training on the test task, however, emergence for MMLU and GSM8K appears to occur at substantially lower scales. We dedicate this section to more closely investigating the relationship between training on the test task and emergence.
@@
We consider only models trained before November 2023, as we have established that these models train on the test task less than newer models. We evaluate the models at intermediate checkpoints as we fine-tune them on the datasets of task relevant data introduced in Section 2.1. We fit $\alpha$ and $c_{e}$ in Equation 1 to the different intermediate checkpoints, and report in Figure 9 top the corresponding points of emergence $c_{e}$. We find that emergence arises at increasingly lower compute regimes as models train on the test task. For MMLU, models exhibit emergence at around $10^{22}$ FLOPs, the scale of Pythia 6.9B. After training on 64,000 examples, emergence arises around $6\cdot 10^{20}$ FLOPs, the scale of Pythia 410M. We observe similar results for GSM8K, see Figure 20 in Appendix E.
Figure 9: Scaling on MMLU as models increasingly train on the test task. The point of emergence $c_{e}$ arises at lower scales (top). Training on the test task yields cleaner log-linear scaling fits (bottom).
@@
The log-linear relationship between pretraining loss and compute is well-established (Kaplan et al., 2020). We observe that training on the test task increasingly recovers log-linear scaling between pretraining compute and benchmark accuracy. Similarly to the earlier section, we evaluate intermediate checkpoints but instead fit log-linear functions in Figure 9 bottom. We observe that the R2 of the fit improves substantially as the models train on more task-relevant data, jumping from $0.63$ to $0.95$ after training on 64,000 examples. Therefore, after training on the test task, almost all the variation in benchmark accuracy is explained by log-linear scaling of pre-training compute. We observe similar results for GSM8K, see Figure 20 in Appendix E.
@@
Schaeffer et al. (2024a) argue that emergence appears due to the choice of metric. To mitigate emergence, they suggest considering Brier score instead of accuracy. We observe, however, that emergence for MMLU does not disappear when using the Brier score, see Figure 5 right, nor that of ARC and HellaSwag when framed as multiple-choice questions, see Figure 19 in Appendix E. We discuss two practical solutions to obtain predictive scaling while maintaining accuracy as the evaluation metric.
For MMLU and multiple-choice benchmarks more broadly, cloze evaluations consistently yield smoother and more predictable scaling, even when using accuracy as the evaluation metric. Since the purpose of these benchmarks is knowledge-testing more so than testing multiple-choice answering ability, cloze evaluations are preferable insofar predictive scaling at lower compute scales is an important consideration. This recommendation aligns with the concurrent work by Gu et al. (2024).
More broadly, if sufficient task relevant data is available, then training on the test task can result in much more predictable scaling by shifting emergence to smaller compute scales. That is, scaling laws where models across scales are fine-tuned on the same, sufficient task-relevant data before evaluation. Such scaling laws correspond to those of specialist models, which for some tasks –e.g., legal annotation (Dominguez-Olmedo et al., 2024)– or purposes –e.g., safety– might be preferable to the scaling laws of generalist models.